OpenStudy (eliesaab):

This a variation of some problem that was proposed earlier. Let \(an\) be a sequence of real numbers defined as follows: \[ a_1=3\\ a_{n+1}=\sqrt{5 a_n-6} \] Show that \(a_n \ge 2.5 \) for every \(n\). Show that \(a_n\) is decreasing. Find its limit.

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle a_1=3 }\) \(\color{black}{\displaystyle a_2=\sqrt{5a_1-6}=\sqrt{9}=3 }\) So for \(n=1\) and \(n=2\), \(a_n=3\). Assume \(\color{black}{\displaystyle a_n=3 }\) (for all, blah blah blah) \(\color{black}{\displaystyle a_{n+1}=\sqrt{5a_n-6} }\) \(\color{black}{\displaystyle a_{n+1}=\sqrt{5\times 3-6}=\sqrt{9}=3 }\)

OpenStudy (solomonzelman):

So you can show by induction ((in fact you don't even need to show the bases for n=2)), that every \(n\), \(a_n=3\), and therefore \(a_n>2.5\).

OpenStudy (solomonzelman):

I would doubt though that \(a_n\) is decreasing, knowing that all terms are the same. Also, having shown \(\forall n\in\mathbb{N}\)\((a_n=3)\), we have the limit of the sequence that way.

OpenStudy (eliesaab):

Sorry, \(a_1=4\)

OpenStudy (solomonzelman):

Oh, that changes everything xD

OpenStudy (solomonzelman):

just informally, \(\color{black}{\displaystyle a_1=4 \ge2.5}\) \(\color{black}{\displaystyle a_2=\sqrt{5(4)-6}=\sqrt{14}\ge3^2 \ge2.5}\) Assume \(\color{black}{\displaystyle a_n\ge2.5}\) blah blah blah \(\color{black}{\displaystyle a_{n+1}\ge\sqrt{5(2.5)-6}\ge\sqrt{5(2.5)-(2.5)(2.5)}\ge2.5}\)

OpenStudy (solomonzelman):

(Knowing that 2.5^2=6.25>6 )

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle a_1=4}\) \(\color{black}{\displaystyle a_2=\sqrt{5(4)-6}=\sqrt{14}<3.8}\) So it is decreasing for the first two terms. Assume \(\color{black}{\displaystyle a_{n-1}>a_{n}}\) for all integers \(1\le k\le n\). Since \(\color{black}{\displaystyle a_{n}=\sqrt{5a_{n-1}-6}}\) (assuming terms are real and defined), therefore \(\color{black}{\displaystyle a_{n}\ge 0}\), which allows, \(\color{black}{\displaystyle \left[a_{n-1}>a_{n}\right]~\Longleftrightarrow \quad (a_{n-1})^2>(a_{n})^2}\). Then, \(\color{black}{\displaystyle (a_{n-1})^2>(a_{n})^2}\) \(\color{black}{\displaystyle (a_{n-1})^2-(a_{n})^2>0}\) \(\color{black}{\displaystyle 5a_n-6-(5a_{n+1}-6)>0}\) \(\color{black}{\displaystyle 5a_n-5a_{n+1}>0}\) \(\color{black}{\displaystyle a_n>a_{n+1}}\).

OpenStudy (solomonzelman):

(Also by induction) \(a_{n+1}<a_{n}\).

OpenStudy (solomonzelman):

If this limit exists, (if sequence converges to some \(x\)) \(x=\sqrt{5x-6}\) \(x^2=5x-6\) \(x^2-5x+6=0\) \((x-2)(x-3)=0\) \(x=2,3\). so, if we prove that \(a_n>2.5\) then we will have the limit, but I am out of time right now. (Shabbat is starting)

OpenStudy (solomonzelman):

I didn't realize but I already proved that \(a_n\ge 2.5\) for all \(n\). What can do, also is to show \(a_n>3\). \(\color{black}{\displaystyle a_1=4>3}\) \(\color{black}{\displaystyle a_2=\sqrt{5(4)-6}=\sqrt{14}>3}\) Assume \(\color{black}{\displaystyle a_n>3}\) ... then, \(\color{black}{\displaystyle a_n>3}\) \(\color{black}{\displaystyle (a_n)^2>9}\) \(\color{black}{\displaystyle (5a_{n+1}-6)>9}\) \(\color{black}{\displaystyle 5a_{n+1}>15}\) \(\color{black}{\displaystyle a_{n+1}>3}\).